OpenStudy (anonymous):

What is the graph of the absolute value equation? y = 5|x|

terenzreignz (terenzreignz):

This is the same as y = |5x| right?

OpenStudy (anonymous):

idk

terenzreignz (terenzreignz):

A nifty little property of absolute values |ab| = |a||b| Since y = 5|x|=|5||x| then y = |5x|

OpenStudy (anonymous):

ok but wht do i do when its like that cuz i know that y=lxl+5 means it gets translated up or that y=lxl-5 gets translated down n that y=lx+5l gets translated to the left etc..

terenzreignz (terenzreignz):

Well, if adding to the absolute value translates the graph multiplying to the absolute value "folds" the graph You might notice that the line becomes steeper the bigger the number you multiply to x, right?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so tht means if the lin is at (0,0) then y=5lxl is 5,5

OpenStudy (anonymous):

wait im confused

OpenStudy (anonymous):

these are the graphs

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Sorry, I got carried away. For any function y = c|x| If c is positive, then the graph should look like a V, which becomes increasingly "folded" the larger c is. If c is negative, then the graph shoud look like a ^, which becomes increasingly "folded" the larger the absolute value of c is... given that, which of your graphs look like V's ?

OpenStudy (anonymous):

the first 2

OpenStudy (anonymous):

one is wider than the other